use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
2x^2 + x - 1 = 0 Answer by venugopalramana(3286) 
 use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
2x^2 + x - 1 = 0...EQN. IS AX^2+BX+C=0...A=2..B=1...C=-1
DISCRIMINANT = B^2-4AC.....
1^2-4*2*(-1)=1+8=9=POSITIVE AND PERFECT SQUARE
HENCE ROOTS ARE REAL AND DISTINCT AND RATIONAL